OpenStudy (anonymous):

Solve?(: x^2+8x-48=0

OpenStudy (anonymous):

-b +- sqrt(b^2 - 4ac) / 2a 8 +- sqrt(64 - 4(1)(-48)) / 2(1) 8 +- sqrt(64 + 192) /2 8 +- sqrt(256) /2 8+- 16 /2 24/2 and -8/2 = 12 and -4 Hope that helps.

OpenStudy (anonymous):

take 48 to lefthand side. Then add 16 to both sides and solve.

OpenStudy (callisto):

Factor the expression on the left. x^2+8x-48=0 x^2 + 12x - 4x - 48 = 0 x ( x+12) - 4 ( x+12) = 0 (x-4)(x+12) = 0 (x-4) = 0 or (x+12) = 0 <- solve these two equations.

OpenStudy (anonymous):

Those are ahrd equations.. i'll try.. Thanks! (:

OpenStudy (anonymous):

I got.. -12 and 4?

OpenStudy (callisto):

That should be correct.

OpenStudy (anonymous):

Thanks so much :D

OpenStudy (callisto):

Welcome. :)